OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

Just foil everything out.

jimthompson5910 (jim_thompson5910):

(x + 4)(x − 4) = x^2-16 is a difference of squares while (x+4)(x+4) = x^2 + 8x + 16 is a perfect square trinomial

jimthompson5910 (jim_thompson5910):

They both share an x^2 term, but everything else is different.